OpenStudy (anonymous):

sin^3 x + cos^3 x over sinx + cosx =1-(sinx)(cosx) verify the Identities

OpenStudy (anonymous):

well..this was asked before,but i can't find the question..so here goes: \[a ^{3} + b ^{3} = (a+b)( a ^{2} -ab +b ^{2}\]

OpenStudy (anonymous):

similiarly..\[\sin ^{3}x + \cos ^{3}x = (sinx +cosx)*(\sin ^{2}x -sinxcosx +\cos ^{2}\]

OpenStudy (anonymous):

oops ... i forgot the closing brackets for both :) you follow @Gabysolis49

OpenStudy (anonymous):

I'm lost^

OpenStudy (anonymous):

its just another way of writing it. if you simplify the rhs..you'l get the same thing okay?

OpenStudy (aivantettet26):

\[\sin^3 x + \cos^3 x \over sinx + cosx\]

OpenStudy (anonymous):

the hard part about this question is getting the idea to write the above.its all downhill from here :)

OpenStudy (aivantettet26):

\[\sin3x+\cos3xsinx+cosx\] = \[1-(sinx)(cosx)\]

OpenStudy (aivantettet26):

i cant seem to draaw the equation

OpenStudy (anonymous):

Yea that's the equation

OpenStudy (aivantettet26):

\[\sin^3 x + \cos^3 x \over sinx + cosx\] \[=1-(sinx)(cosx) \]

OpenStudy (anonymous):

hey you guys..listen first

OpenStudy (anonymous):

@aivantettet26 and @Gabysolis49 do you get how i changed the numerator??

OpenStudy (aivantettet26):

we can divide sin^3x + cos^3x / sinx + cos x

OpenStudy (aivantettet26):

okay? im listening XD

OpenStudy (anonymous):

right check how i changed the numerator

OpenStudy (aivantettet26):

okay okay. im following

OpenStudy (anonymous):

Okay I just need help answering it ^^^

OpenStudy (anonymous):

right here goes:

OpenStudy (anonymous):

???^

OpenStudy (anonymous):

\ (sinx +cosx)*(\sin ^{2}x -sinxcosx+\cos ^{2}x)\div(sinx +cosx)\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\[(sinx +cosx)*(\sin ^{2}x -sinxcosx+\cos ^{2}x)\div(sinx +cosx)\]

OpenStudy (anonymous):

sorry guys..connection problems

OpenStudy (anonymous):

@Gabysolis49 @aivantettet26

OpenStudy (anonymous):

Okay continue please

OpenStudy (anonymous):

i want u to do it now

OpenStudy (anonymous):

see if you can cancel anything

OpenStudy (anonymous):

Yea sinx and cosx

OpenStudy (anonymous):

okay so write the new equation:

OpenStudy (anonymous):

(Sinx+cosx)(sin^2 x+cos^2 x)

OpenStudy (anonymous):

sorry..thats not it..try again :)

OpenStudy (anonymous):

Sin^2 x-sinx cosx+cos^2 x

OpenStudy (anonymous):

yah!

OpenStudy (anonymous):

So how does that equal 1-sinx cosx

OpenStudy (anonymous):

hmmm what is 1??? :)

OpenStudy (anonymous):

@Gabysolis49

OpenStudy (anonymous):

sin3x+cos3xsinx+cosx =1−(sinx)(cosx)

OpenStudy (anonymous):

sin^2x+cos^2x=1. you got it now ?